A positive number is 5 times another number. If 21 is added to both the numbers, then one of the new numbers becomes twice the other new number. What are the numbers? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let one of the positive number be x then other number will be 5x.
According to the question,
5x+21=2(x+21)
⇒5x+21=2x+42
⇒5x-2x=42-21
⇒3x=21
⇒x=7
Onenumber=x=7
Other number = 5x = 5×7 = 35
The two numbers are 7 and 35.

Thanks